OpenStudy (anonymous):

Judah makes pottery to sell at his school’s craft bazaar. Every year, he determined that he makes fifteen dollars less than he made the previous year. Write a function rule to define the total amount y Judah made based on the amount x he made the previous year. Give an example of how the function rule works. Which is the dependent variable? The independent variable?

OpenStudy (anonymous):

well what do you think

OpenStudy (anonymous):

x + y = 15?

OpenStudy (anonymous):

he makes 15 LESS every year. so this year = last year - 15. what is this when substituting for variables>

OpenStudy (anonymous):

y = x + b y + - 15 + b The dependent variable would be y.

OpenStudy (anonymous):

I mean y= -15 + b